# Thera: Aliasing-Free Arbitrary-Scale Super-Resolution with Neural Heat Fields

### v3

@@ Abstract
1 Introduction
Over the years, learning-based image super-resolution (SR) methods have achieved increasingly better results. However, unlike interpolation techniques that can resample images at any resolution, these methods typically require retraining for each scaling factor. Recently, arbitrary-scale SR (ASR) approaches have emerged, which allow users to specify any desired scaling factor without retraining, significantly increasing flexibility [18]. Notably, with LIIF, Chen et al. [10] pioneered the use of neural fields for single-image SR, exploiting their continuous representation to enable SR at arbitrary scaling factors. LIIF has since inspired several follow-ups which build upon the idea of using per-pixel neural fields [24, 7, 9, 56]. This is not surprising: Neural fields are in many ways a natural match for variable-resolution computer vision and graphics [48]. By implicitly parameterizing a target signal as a neural network that maps coordinates to signal value, they offer a compact representation, defined over a continuous input domain, and are analytically differentiable.
@@ 1 Introduction
While neural fields naturally model continuous functions, they do not easily allow for observations of such functions other than point-wise evaluations. For many tasks, however, integral observation models such as point spread functions (PSFs) are desirable. This is particularly true for neural fields-based ASR methods, which by nature do not commit to a fixed upscaling factor a priori but regress continuous representations with unbounded spectra that can be observed at various sampling rates. If the Nyquist frequency corresponding to the desired sampling rate is lower than the highest frequency represented by the field, the sampling operation is prone to aliasing. This explains the initially counterintuitive relevance of anti-aliasing for super-resolution: When using neural fields, signals are first upsampled to infinite (continuous) resolution and then resampled at the desired resolution, and this latter operation must be done carefully. Incorporating a physically plausible observation model is not trivial [2, 3, 27, 51, 17, 4], but has the potential to avoid aliasing. For this reason, Chen et al. [10] and successor works [24, 7, 9, 56] have already taken a first step towards learning multi-scale representations, via cell encoding. Fundamentally, these “learning-based anti-aliasing” approaches try to learn an integrated version of the field for any scaling factor, wasting field capacity for approximating a relation that can be described exactly through Fourier theory.
@@
$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$31.531.531.531.531.631.631.631.631.731.731.731.731.831.831.831.831.931.931.931.93232323232.132.132.132.1Thera AirThera PlusThera Pro(Ours)MetaSRLIIFCUFCiaoSRCLITMSITSRNOModel Size (M parameters)PSNR (dB)
@@
Figure 2: Comparison of recent ASR methods, averaged over ×$\times\{2,3,4\}$ scales. We generally achieve higher performance at lower parameter counts. Our best model, Thera Pro, achieves highest overall performance by a large margin.
@@
In this work, we combine recent advances in implicit neural representations with ideas from classical signal theory to introduce neural heat fields, a novel type of neural field that guarantees anti-aliasing by construction. The key insight is that sinusoidal activation functions [41] enable selective attenuation of individual components depending on their spatial frequency, following Fourier theory. This allows for the exact computation of Gaussian-blurred versions of the field for any desired (isotropic) blur radius. When rasterizing an image, the field can therefore be queried with a Gaussian PSF that matches the target resolution, effectively preventing aliasing. Notably, filtering with neural heat fields incurs no computational overhead: The querying cost is the same for any width of the anti-aliasing filter kernel, including infinite and zero widths.
@@
Building on this, we then propose Thera, an end-to-end ASR method that combines a hypernetwork [16] with a grid of local neural heat fields, offering theoretical guarantees with respect to multi-scale representation (see Fig. 1). Empirically, Thera outperforms all competing ASR methods, often by a substantial margin, and is more parameter-efficient (see Fig. 2). To the best of our knowledge, Thera is also the first neural field method to allow bandwidth control at test time.
@@ 2.1 Neural Fields
A neural field, also called an implicit neural representation, is a neural network trained to map coordinates onto values of some physical quantity. Recently, neural fields have been used for parameterizing various types of visual data, including images [22, 41, 42, 10, 24, 12, 47], 3D scenes (e.g., represented as signed distance fields [36, 41, 40, 46, 47], occupancy fields [33, 37], LiDAR fields [20], view-dependent radiance fields [34, 2, 3, 4, 47]), or digital humans [52, 55, 8, 49, 15]. Frequently, it is desirable to impose some prior over the space of learnable implicit representations. A common approach for such conditioning is encoder-based inference [48], where a parametric encoder maps input observations to a set of latent codes $\bm{z}$, which are often local [10, 24, 44, 7, 9]. The encoded latent variables $\bm{z}$ are then used to condition the neural field, for instance by concatenating $\bm{z}$ to the coordinate inputs or through a more expressive hypernetwork [16], mapping latent codes $\bm{z}$ to neural field parameters $\bm{\theta}$. An early example of this approach, which is gaining popularity [48], was proposed in [41].
@@ 2.2 Arbitrary-Scale Super-Resolution
ASR is the sub-field of single-image SR in which the desired SR scaling factor can be chosen at inference time to be (theoretically) any positive number, allowing maximum flexibility, such as that of interpolation methods. The first work along this line is MetaSR [18], which infers the parameters of a convolutional upsampling layer using a hypernetwork [16] conditioned on the desired scaling factor. An influential successor work is LIIF [10], in which the high-resolution image is implicitly described by local neural fields. These fields are conditioned via concatenation, with features extracted from the low-resolution input image. The continuous nature of the neural fields allows for sampling target pixels at arbitrary locations and thus also arbitrary resolution.
Most subsequent work has since been built upon the LIIF framework. For example, UltraSR [50] improves the modeling of high-frequency textures with periodic positional encodings of the coordinate space, as is common practice for e.g., neural radiance fields [34, 2, 3, 4]. LTE [24] makes learning higher frequencies more explicit by effectively implementing a learnable coordinate transformation into 2D Fourier space, prior to a forward pass through an MLP. Vasconcelos et al. [44] use neural fields in CUF to parameterize continuous upsampling filters, which enables arbitrary-scale upsampling. More recently, methods like CiaoSR [7], CLIT [9], and most recently MSIT [56] have integrated (multi-scale) attention mechanisms, improving reconstruction quality. In a parallel line of research, Wei et al. [45] propose SRNO, an attention-based neural operator that learns a continuous mapping between low- and high-resolution function spaces.
@@ 2.3 Anti-Aliasing in Neural Fields
Early in the recent development of implicit neural representations, concerns regarding aliasing were raised. Barron et al. [2] proposed integrating a positional encoding with Gaussian weights, which reduced aliasing in NeRF [34]. Improvements were later proposed for unbounded scenes [3] and to improve efficiency [17]. In [4], Barron et al. tackle anti-aliasing within the Instant-NGP [35] approach. Recent work has succeeded in limiting the bandwidth using multiplicative filter networks [27], polynomial neural fields [51] or cascaded training [39], although these works are restricted to discrete, pre-defined band limits (and thus resolutions) and have not tackled super-resolution tasks. These methods are not a good fit for ASR because they do not allow for continuous anti-aliasing, nor bandwidth control at test time. To perform scale-dependent filtering, most fields-based ASR methods instead explicitly provide the scale as input to the field, attempting to learn an appropriate observation model from data. While this approach may work reasonably well in in-distribution settings, it seeks to learn a model from data that can be described exactly with a differential equation, ultimately sacrificing fidelity and generalization.
@@ 3.1 Neural Heat Fields for Analytical Anti-Aliasing
Let $\mathbf{x}\in\mathbb{R}^{2}$ denote the spatial coordinates of a continuous image function $f(\mathbf{x})$. Aliasing occurs when this continuous signal is sampled at a rate that does not adequately capture its highest frequency components, resulting in overlapping spectral replicas in the Fourier domain. One must therefore apply a low-pass filter $g(\mathbf{x})$ whose cut-off frequency is aligned with the Nyquist frequency of the sampling rate, then sample the band-limited signal $f\circledast g(\mathbf{x})$. The key of our method is that, if a signal is decomposed into sinusoidal components, such filtering can be done simply by re-scaling each component by a factor that depends on their frequency and a time coordinate $t$. This behavior is naturally accomplished by parameterizing the field $\Phi$ as a two-layer perceptron,
@@
Here, $\left|\nu\right|=\left|\nu(\mathbf{W}_{1})\right|$ denotes the row-wise Euclidean norm of $\mathbf{W}_{1}$, representing the magnitudes of the implied wave numbers (frequencies). Interestingly, Eq. 1 constitutes the solution of the isotropic heat equation $\frac{\partial\Phi}{\partial t}=\kappa\cdot\nabla_{\mathbf{x}}^{2}\Phi$, as derived in Appendix A. We therefore refer to this MLP as a neural heat field.
There is an ideal bijection between the desired sampling rate $f_{s}$ and $t$. At $t=0$ no filtering takes place, implying a continuous signal ($f_{s}\to\infty$). A low-pass filtered version of the signal is observed for $t>0$. To obtain a desired level of anti-aliasing, we only need to compute the corresponding value of $t$. The relationship between the cut-off frequency of the filter and $t$ is controlled by the diffusivity constant $\kappa$, which we can freely set to any positive number. For simplicity, and without loss of generality, we set $\kappa$ such that the native resolution of the (observed, discrete) signal $\mathcal{D}$ is $t=1$. Assuming equal sampling rate $f_{s}$ along both coordinate axes, the optimal value of $\kappa$ then evaluates to
$\kappa=\frac{\sqrt{\ln(4)}}{2f_{s}\pi^{2}}.$ (3)
@@
$t=S^{2}.$ (4)
For a derivation of these values and a demonstration of the filtering mechanism of neural heat fields, see Appendix A.
@@ 3.2 Learning a Super-Resolution Prior
The multi-scale signal representation inherent in neural heat fields is a natural match for ASR. Still, two challenges must be addressed. First, our formulation restricts the choice of architecture to MLPs with a single hidden layer. Second, while it theoretically guarantees the downsampling operation ($t>1$), the upsampling operation ($10<t<$ 1) remains ill-posed. To narrow down the (infinite) solution space to a unique result, a prior must either be defined in an unsupervised fashion or learned from data. Our solution to both challenges is to condition local fields with a hypernetwork $\Psi:\mathbb{R}^{W\times H\times 3}\to\mathbb{R}^{W\times H\times N}$, mapping the low-resolution input image to the $N$ parameters of each local neural heat field. As originally proposed by LIIF [10] and adopted by recent ASR methods [24, 7, 44, 9], each local field spans the area of one pixel of the low-resolution input. It is important that even though the fields themselves model only a local part of the image, the hypernetwork informs them with contextual features collected over a large receptive field. The hypernetwork $\Psi$ is composed of a standard SR backbone, as used in previous work [10, 24, 44, 7, 9], followed by one or more layers that produce feature maps of depth $N$.
@@
The described scheme, which we call Thera, is depicted in Fig. 3. It allows for arbitrary-scale super-resolution, combining the multi-scale signal representation within neural heat fields with the expressivity of proven feature extraction backbones for SR and image restoration. As the entire network is trained end-to-end, the feature extractor can learn super-resolution priors for a whole range of resolutions covered by the training data. E.g., a network trained with scaling factors up to $\times$4 will encode priors that enable us to observe the field at $t>\frac{1}{16}$. By training on multiple resolutions, we can also make $\kappa$ a trainable parameter that allows the network to adapt to different downsampling operators. Finally, we set the bias terms for the three color channels of every local field $\Phi$ (i.e., $\textbf{b}_{2}$) to the RGB values of the associated low-resolution pixel. Thus, the hypernetwork only predicts field-wise phase shifts $\mathbf{b}_{1}$ and amplitudes $\mathbf{W}_{2}$.

### v4

@@ Abstract
1 Introduction
Over the years, learning-based image super-resolution (SR) methods have achieved increasingly better results. However, unlike interpolation techniques that can resample images at any resolution, these methods typically require retraining for each scaling factor. Recently, arbitrary-scale SR (ASR) approaches have emerged, which allow users to specify any desired scaling factor without retraining, significantly increasing flexibility (Hu et al., 2019). Notably, with LIIF, Chen et al. (2021) pioneered the use of neural fields for single-image SR, exploiting their continuous representation to enable SR at arbitrary scaling factors. LIIF has since inspired several follow-ups which build upon the idea of using per-pixel neural fields (Lee & Jin, 2022; Cao et al., 2023; Chen et al., 2023; Zhu et al., 2025). This is not surprising: Neural fields are in many ways a natural match for variable-resolution computer vision and graphics (Xie et al., 2022). By implicitly parameterizing a target signal as a neural network that maps coordinates to signal value, they offer a compact representation, defined over a continuous input domain, and are analytically differentiable.
@@ 1 Introduction
While neural fields naturally model continuous functions, they do not easily allow for observations of such functions other than point-wise evaluations. For many tasks, however, integral observation models such as point spread functions (PSFs) are desirable. This is particularly true for neural fields-based ASR methods, which by nature do not commit to a fixed upscaling factor a priori but regress continuous representations with unbounded spectra that can be observed at various sampling rates. If the Nyquist frequency corresponding to the desired sampling rate is lower than the highest frequency represented by the field, the sampling operation is prone to aliasing. This explains the initially counterintuitive relevance of anti-aliasing for super-resolution: When using neural fields, signals are first upsampled to infinite (continuous) resolution and then resampled at the desired resolution, and this latter operation must be done carefully. Incorporating a physically plausible observation model is not trivial (Barron et al., 2021; 2022; Lindell et al., 2022; Yang et al., 2022; Hu et al., 2023; Barron et al., 2023), but has the potential to avoid aliasing. For this reason, Chen et al. (2021) and successor works (Lee & Jin, 2022; Cao et al., 2023; Chen et al., 2023; Zhu et al., 2025) have already taken a first step towards learning multi-scale representations, via cell encoding. Fundamentally, these “learning-based anti-aliasing” approaches require the scaling factor (or, equivalently, the output pixel area) as additional input to the neural field and learn an integrated (i.e., appropriately blurred and therefore anti-aliased) version of the field for each scaling factor; arguably wasting field capacity to approximate a relation that can be described exactly through Fourier theory.
@@
$10^{-2}$$^{-1}1001$$^{0}$$10^{1}$31.531.531.631.631.731.731.831.831.931.9323232.132.1Thera AirThera PlusThera Pro(Ours)MetaSRLIIFCUFCiaoSRCLITMSITSRNOModel Size (M parameters)PSNR (dB)
@@
Figure 2: Comparison of recent ASR methods, averaged over $\times\{2,3,4\}$ scales. We generally achieve higher performance at lower parameter counts. Our best model, Thera Pro, achieves highest overall performance by a large margin.
@@
In this work, we combine recent advances in implicit neural representations with ideas from classical signal theory to introduce neural heat fields, a novel type of neural field that guarantees anti-aliasing by construction. The key insight is that sinusoidal activation functions (Sitzmann et al., 2020b) enable selective attenuation of individual components depending on their spatial frequency, following Fourier theory. This allows for the exact computation of Gaussian-blurred versions of the field for any desired (isotropic) blur radius. When rasterizing an image, the field can therefore be queried with a Gaussian PSF that matches the target resolution, effectively preventing aliasing. In practice, heat fields receive an additional input coordinate $t$, controlling the strength of the Gaussian blur applied to the signal. Unlike learning-based anti-aliasing, the resulting filtering operation is expressed analytically, rather than learned from data. In other words, previous approaches fit a 3D field ($x$, $y$, and scalescale) while we only need to fit a 2D field ($x$ and $y$), whereas the scale dimension is computed analytically, significantly reducing field complexity and data requirements. Notably, filtering with neural heat fields incurs no computational overhead: The querying cost is the same for any width of the anti-aliasing filter kernel, including infinite and zero widths.
@@
Building on this, we then propose Thera, an end-to-end ASR method that combines a hypernetwork (Ha et al., 2017) with a grid of local neural heat fields, offering theoretical guarantees with respect to multi-scale representation (see Figure 1). Empirically, Thera outperforms all competing ASR methods, often by a substantial margin, and is more parameter-efficient (see Figure 2). To the best of our knowledge, Thera is also the first neural field method to allow bandwidth control at test time.
@@ 2.1 Neural Fields
A neural field, also called an implicit neural representation, is a neural network trained to map coordinates onto values of some physical quantity. Recently, neural fields have been used for parameterizing various types of visual data, including images (Karras et al., 2021; Sitzmann et al., 2020b; Tancik et al., 2020; Chen et al., 2021; Lee & Jin, 2022; de Lutio et al., 2019; Wu et al., 2023), 3D scenes (e.g., represented as signed distance fields (Park et al., 2019; Sitzmann et al., 2020b; a; Williams et al., 2022; Wu et al., 2023), occupancy fields (Mescheder et al., 2019; Peng et al., 2020), LiDAR fields (Huang et al., 2023), view-dependent radiance fields (Mildenhall et al., 2021; Barron et al., 2021; 2022; 2023; Wu et al., 2023)), or digital humans (Yenamandra et al., 2021; Zheng et al., 2022; Cao et al., 2022; Xiu et al., 2022; Giebenhain et al., 2023). Frequently, it is desirable to impose some prior over the space of learnable implicit representations. A common approach for such conditioning is encoder-based inference (Xie et al., 2022), where a parametric encoder maps input observations to a set of latent codes $\bm{z}$, which are often local (Chen et al., 2021; Lee & Jin, 2022; Vasconcelos et al., 2023; Cao et al., 2023; Chen et al., 2023). The encoded latent variables $\bm{z}$ are then used to condition the neural field, for instance by concatenating $\bm{z}$ to the coordinate inputs or through a more expressive hypernetwork (Ha et al., 2017), mapping latent codes $\bm{z}$ to neural field parameters $\bm{\theta}$. An early example of this approach, which is gaining popularity (Xie et al., 2022), was proposed in Sitzmann et al. (2020b).
@@ 2.2 Arbitrary-Scale Super-Resolution
ASR is the sub-field of single-image SR in which the desired SR scaling factor can be chosen at inference time to be (theoretically) any positive number, allowing maximum flexibility, such as that of interpolation methods. The first work along this line is MetaSR (Hu et al., 2019), which infers the parameters of a convolutional upsampling layer using a hypernetwork conditioned on the desired scaling factor. An influential successor work is LIIF (Chen et al., 2021), in which the high-resolution image is implicitly described by local neural fields. These fields are conditioned via concatenation, with features extracted from the low-resolution input image. The continuous nature of the neural fields allows for sampling target pixels at arbitrary locations and thus also arbitrary resolution.
Most subsequent work has since been built upon the LIIF framework. For example, UltraSR (Xu et al., 2021) improves the modeling of high-frequency textures with periodic positional encodings of the coordinate space, as is common practice for e.g., neural radiance fields (Mildenhall et al., 2021; Barron et al., 2021; 2022; 2023). LTE (Lee & Jin, 2022) makes learning higher frequencies more explicit by effectively implementing a learnable coordinate transformation into 2D Fourier space, prior to a forward pass through an MLP. Vasconcelos et al. (2023) use neural fields in CUF to parameterize continuous upsampling filters, which enables arbitrary-scale upsampling. More recently, methods like CiaoSR (Cao et al., 2023), CLIT (Chen et al., 2023), and most recently MSIT (Zhu et al., 2025) have integrated (multi-scale) attention mechanisms, improving reconstruction quality. In a parallel line of research, Wei & Zhang (2023) propose SRNO, an attention-based neural operator that learns a continuous mapping between low- and high-resolution function spaces.
Another line of work employs generative models such as denoising diffusion for SR (Saharia et al., 2022; Gao et al., 2023). While most methods minimize per-pixel errors (essentially predicting the minimum mean square error estimate), generative models are trained to produce more realistic-looking outputs by predicting one of many plausible high-resolution images. However, since specific ground truth details are not exactly recovered, such models typically report worse distortion metrics (like PSNR and SSIM) compared to pixel-based methods, c.f. Blau & Michaeli (2018); Delbracio & Milanfar (2023). In this paper, we adopt a pixel-based objective to preserve fidelity to the ground truth, which is important for many downstream applications (e.g., face or license plate recognition).
@@ 2.3 Anti-Aliasing in Neural Fields
Early in the recent development of implicit neural representations, concerns regarding aliasing were raised. Barron et al. (2021) proposed integrating a positional encoding with Gaussian weights, which reduced aliasing in NeRF (Mildenhall et al., 2021). Improvements were later proposed for unbounded scenes (Barron et al., 2022) and to improve efficiency (Hu et al., 2023). Barron et al. (2023) tackle anti-aliasing within the Instant-NGP (Müller et al., 2022) approach. Recent work has succeeded in limiting the bandwidth using multiplicative filter networks (Lindell et al., 2022), polynomial neural fields (Yang et al., 2022) or cascaded training (Shabanov et al., 2024), although these works are restricted to discrete, pre-defined band limits (and thus resolutions) and have not tackled super-resolution tasks. These methods are not a good fit for ASR because they do not allow for continuous anti-aliasing, nor bandwidth control at test time. To perform scale-dependent filtering, most fields-based ASR methods instead explicitly provide the scale as input to the field, attempting to learn an appropriate observation model from data. While this approach may work reasonably well in in-distribution settings, it seeks to learn a model from data that can be described exactly with a differential equation, ultimately sacrificing fidelity and generalization.
@@ 3.1 Neural Heat Fields for Analytical Anti-Aliasing
Let $\mathbf{x}\in\mathbb{R}^{2}$ denote the spatial coordinates of a continuous image function $f(\mathbf{x})$. Aliasing occurs when this continuous signal is sampled at a rate that does not adequately capture its highest frequency components, resulting in overlapping spectral replicas in the Fourier domain. One must therefore apply a low-pass filter $g(\mathbf{x})$ whose cut-off frequency is aligned with the Nyquist frequency of the sampling rate, then sample the band-limited signal $f\circledast g(\mathbf{x})$. The key of our method is that, if a signal is decomposed into sinusoidal components, such filtering can be done simply by re-scaling each component by a factor that depends on their frequency as well as a time coordinate, which we call $t$. The time coordinate acts as a third, continuous input to the neural field and controls the amount of re-scaling, and therefore the Gaussian blur applied to the signal. This perfectly mimics how high-frequency components decay faster than low-frequency ones in the analytical solution to the heat equation. The detailed derivation can be found in Appendix A. The behavior described above is naturally accomplished by parameterizing the field $\Phi$ as a two-layer perceptron,
@@
Here, $\left|\nu\right|=\left|\nu(\mathbf{W}_{1})\right|$ denotes the row-wise Euclidean norm of $\mathbf{W}_{1}$, representing the magnitudes of the implied wave numbers (frequencies). Interestingly, Equation 1 constitutes the solution of the isotropic heat equation $\frac{\partial\Phi}{\partial t}=\kappa\cdot\nabla_{\mathbf{x}}^{2}\Phi$, as derived in Appendix A. We therefore refer to this MLP as a neural heat field.
There is an ideal bijection between the desired sampling rate $f_{s}$ and $t$. At $t=0$ no filtering takes place, implying a continuous signal ($f_{s}\to\infty$). A low-pass filtered version of the signal is observed for $t>0$. To obtain a desired level of anti-aliasing, we only need to compute the corresponding value of $t$. The relationship between the cut-off frequency of the filter and $t$ is controlled by a global diffusivity constant $\kappa$, which defines how fast components of different frequencies decay over time in the underlying PDE model. We can freely set $\kappa$ to any positive number, but for simplicity, and without loss of generality, we set $\kappa$ in our theoretical derivations such that the native resolution of the (observed, discrete) signal $\mathcal{D}$ corresponds to $t=1$. Assuming equal sampling rate $f_{s}$ along both coordinate axes, the optimal value of $\kappa$ then evaluates to
$\kappa=\frac{\ln(4)}{2f_{s}^{2}\pi^{2}}.$ (3)
@@
$t=S^{2},$ (4)
where $S$ is the subsampling rate, i.e., the inverse of the scaling factor. In other words, the equation above defines the correct value for $t$ according to the scaling factor at which the field should be sampled. For a derivation of these values and a demonstration of the filtering mechanism of neural heat fields, see Section A.
@@ 3.2 Learning a Super-Resolution Prior
The multi-scale signal representation inherent in neural heat fields is a natural match for ASR. Still, two challenges must be addressed. First, our formulation restricts the choice of architecture to MLPs with a single hidden layer. Second, while it theoretically guarantees the downsampling operation ($t>1$), the upsampling operation ($0<t<1$) remains ill-posed. To narrow down the (infinite) solution space to a unique result, a prior must either be defined in an unsupervised fashion or learned from data. Our solution to both challenges is to condition local fields with a hypernetwork $\Psi:\mathbb{R}^{W\times H\times C}\to\mathbb{R}^{W\times H\times N}$. First, a standard backbone, as used in previous work (Chen et al., 2021; Lee & Jin, 2022; Vasconcelos et al., 2023; Cao et al., 2023; Chen et al., 2023), extracts image features from the low-resolution input image. Then, the hypernetwork maps these features to the $N$ parameters of each local neural heat field. As originally proposed by LIIF (Chen et al., 2021) and adopted by recent ASR methods (Lee & Jin, 2022; Cao et al., 2023; Vasconcelos et al., 2023; Chen et al., 2023), each local field spans the area of one pixel of the low-resolution input. It is important that even though the fields themselves model only a local part of the image, the hypernetwork informs them with contextual features collected over a large receptive field.
@@
The described scheme, which we call Thera, is depicted in Figure 3. It allows for arbitrary-scale super-resolution, combining the multi-scale signal representation within neural heat fields with the expressivity of proven feature extraction backbones for SR and image restoration. As the entire network is trained end-to-end, the feature extractor can learn super-resolution priors for a whole range of resolutions covered by the training data. E.g., a network trained with scaling factors up to $\times$4 will encode priors that enable us to observe the field at $t>\frac{1}{16}$. By training on multiple resolutions, we can also make $\kappa$ a trainable parameter that allows the network to adapt to different downsampling operators. Finally, we set the bias terms for the three color channels of every local field $\Phi$ (i.e., $\textbf{b}_{2}$) to the RGB values of the associated low-resolution pixel. Thus, the hypernetwork only predicts field-wise phase shifts $\mathbf{b}_{1}$ and amplitudes $\mathbf{W}_{2}$.